1)
Use the vectors u = 2i - j, v = 21 - 3j, and w = -3i + 5j to evaluate the expression. 2v - u + w
2)

Find a unit vector in the same direction as the given vector. a = 201 - 21j
Answer #1

3 Given ५-२८-j, V=२1-35 and w=-30+55 South २v-u+= २(२८-31)-(२i-j)-31+51 २v-u+ २V-4+w -- & wnit vector in sami disection of a